moonlightinparis07:

A painter made the table below to show the amount of paint it takes to cover a certain amount of square feet. The amount of paint varies directly with the area of each wall. Amount of Paint and Size of Wall Area of wall feet squared, x 64 144 400 Amount of paint (pints), y 4 9 25 How much paint should the painter expect to use to paint a wall with an area of 256 square feet? 16 pints 32 pints 128 pints 225 pints

XioGonz:

There are 2 walls with an area of 20*8 2 walls that are 18*8. Therefore giving us a total area of: 160*2 + 144*2 = 608 square feet The ceiling 20*18 = 360 square feet Total pr c: 968 square feet. Multiply by 2 for 2 c 2*968 = 1936 Divide by 400 The nearest half gallon: 5.0 gallons If you round to the nearest .5 it'll be 4.84

moonlightinparis07:

Never Mind the Answer is 16 pints

jhonyy9:

 so if you check this my posted image - table - you see that the value of x is always equal 16y 64 = 16*4 144 = 16*9 400 = 16*25 so in this way you can calcule easy to get 256 = 16 *x x = ? hope this help you understanding easy @XioGonz @Laylalyssa

XioGonz:

Good Job jhony!

jhonyy9:

ty